# LakeVilla: Multi-Table Transactions for Lakehouses

### v1

@@ 2. The Lakehouse Design Area and Structure
Different Solution Approaches. When addressing missing LH features, we observed four design directions described as questions:
@@
(Q1) What concurrency mechanism should we use? Exclusive locking (Armbrust et al., 2020) prevents concurrent writes while aborting all conflicts. A branching solution, like Project Nessie (licence, 2024), isolates transactions in different branches (Foundation, 2024d). Lastly, presumably Delta Lake’s newly announced commit owner (Jain and Databricks, 2024) applies versioning-based conflict detection and resolution (Armbrust et al., 2020; Foundation, 2024a) known from DB research (Bernstein and Goodman, 1983; Bernstein et al., 1987; Neumann et al., 2015).
@@
(Q2) Do we implement an internal or external solution? Many LH projects (XTable, 2024; licence, 2024; Jain and Databricks, 2024) add external services above the object store. This essentially provides a simple and performant solution. In contrast, a built-in solution within the OTF design without auxiliary architectures reduces dependencies on such services.
@@
(Q3) Should we work with existing OTFs or redefine them? Delta Lake 4.0 update changes its original design and provides backward compatibility with its backfilling feature (Jain and Databricks, 2024). In contrast, adaptive approaches like XTable (XTable, 2024) and Project Nessie (licence, 2024) integrate into existing OTF architectures without noticeable structure changes, providing easier adaption on already deployed systems.
@@
(Q4) For which type of transactions do we use which solution? Most LH projects are one-fit-all solutions affecting all transaction types (Armbrust et al., 2020; Jain and Databricks, 2024; XTable, 2024; licence, 2024). This contradicts a flexible solution, providing individual choice between guarantees. However, this limits its simplicity, as such solutions must offer modular building blocks.
3. LakeVilla
@@
Our solution. We introduce LakeVilla (LV) as a versioning-based (Q1), built-in (Q2), adaptive (Q3), and flexible (Q4) OTF extension providing the missing features (i) – (iii). LV provides three modular features: LV [R] for (i) Recovery, LV [CT] for (ii) complex transactions, and LV [I] for (iii) isolation. Our design can seamlessly integrate into OTFs such as Apache Iceberg and Delta Lake without reformatting or reconfiguring the whole system.
3.1. LakeVilla [Recovery] (LV [R])
@@
Why multi-query transactions abort. When operating with concurrent transactions, the key challenge to prevent aborts is adapting to changes in the table versions of the metadata layer. By default, OTFs apply optimistic concurrency control that checks for conflicts at the commit of a transaction (Projects, 2024b; Foundation, 2024f; Zaharia et al., 2021). This approach assumes that most operations do not conflict. However, OTFs also provide table-level granularity, meaning that all concurrent write operations on the same table will conflict without fine-grained conflict management. Hence, we propose LV [R] to provide recovery (i). This feature combines two techniques we designed for OTFs following our decisions for Q1–Q4: Our markers (technique #1) reserve snapshot IDs at a transaction’s beginning seamlessly and publish a transaction’s changes atomically. To adapt to concurrent changes and faults, we propose our sublog (technique #2) in LV [R].
(a) Apache Iceberg Integration
(b) Delta Lake Integration
@@
Figure 1. Technique #1: Markers (green) reserve snapshot versions in the metadata layers
[…]
[Technique 1: We place file-based markers to reserve snapshot versions.]Technique 1: We place file-based markers in the metadata layer of OTFs to reserve snapshot versions. For Delta Lake, it appends to its log structure. The Iceberg integration places markers on a new path.
@@
Technique #1: Markers We propose file-based markers within OTF metadata layers to prevent transaction starvation. For OTFs, they atomically reserve a table snapshot version during their transactions’ execution, thus blocking concurrent transactions from claiming the same version. We ensure their uniqueness and atomicity by using Put-If-Absent or conditional writes used in OTFs (Amazon Web Services, 2014; Armbrust et al., 2020). Additionally, we implement continuous checks for object stores without atomic requests accessing each marker’s HEAD to detect potential overwrites. If a client fails to claim a version or its marker was deleted by someone else, it aborts or claims a new marker. Markers are temporary objects that must be deleted or overwritten.
Marker integration. Their OTF integration shown in Fig. 1 is not arbitrary. In Iceberg, we place markers in the area of manifest lists and link to the snapshot. For Delta Tables, we append markers to the log directory. We show both approaches in Fig. 1(a) and Fig. 1(b) for marker ‘M1’, respectively. In both integrations, the marker becomes an exclusive lock (Bernstein et al., 2011; Liu et al., 2020; Iwabuchi et al., 2016; Bernstein et al., 1987) to prevent other concurrent transactions from claiming the same snapshot version (i+1).
@@
Marker Compatibility. Following our design principle of flexibility (Q4), we propose markers should not enforce the same guarantees or new techniques to all transactions. Hence, our markers must fulfill their purpose while still being compatible with the original OTF protocols. LV’s markers use the extensible open-source formats (e. g., JSON) of the original logs/manifest and appear empty for the original protocols. At the same time, they are detectable for our marker-based clients with the data contained.
@@
Marker Commit. Nevertheless, commits using markers should be atomically visible. So, we designed a commit phase that replaces these markers noticeable to legacy readers: Our client collects and summarizes all relevant information according to the OTF’s required instructions during a commit. If needed, it builds underlying structures, such as Iceberg’s manifest lists, and finally overwrites the placed marker with all the necessary instructions. This action effectively makes the whole transaction result of a table available even for the original OTF protocols. In case of an abort, our client overwrites a marker with a previous manifest list or an empty log entry to not disturb the original protocols.

### v2

@@ 3.2 Design Questions
When addressing missing LH features and exploring prior approaches (e.g., delta4_multi_table ; ducklake ), we observed four design directions described as questions:
@@
(Q1) What concurrency mechanism should we use? Exclusive locking journals/pvldb/ArmbrustDPXZ0YM20 prevents concurrent writes while aborting all conflicts. A branching solution, like Project Nessie project_nessie_web , isolates transactions in different branches iceberg_nessie_doc . Lastly, Delta Lake’s newly announced commit owner delta4_multi_table and DuckLake ducklake apply versioning-based conflict detection and resolution journals/pvldb/ArmbrustDPXZ0YM20 ; Iceberg_website known from DB research journals/tods/BernsteinG83 ; books/aw/BernsteinHG87 ; conf/sigmod/0001MK15 .
@@
(Q2) Do we implement an internal or external solution? Most LH projects extract responsibilities from the OTF to other external services above the object store delta4_multi_table ; project_nessie_web ; ducklake ; xtable_web or to their database ducklake . This essentially provides a simple and performant solution. In contrast, a built-in solution within the OTF design without auxiliary architectures reduces dependencies on such services.
@@
(Q3) Should we work with existing OTFs or redefine them? Delta Lake 4.0 updates its original design and provides backward compatibility with its backfilling feature delta4_multi_table . Further, DuckLake manages metadata exclusively within their database, such that only the raw data files remain on the object store ducklake . In contrast, adaptive approaches like XTable xtable_web and Project Nessie project_nessie_web integrate into existing OTF architectures without noticeable structure changes, providing easier adoption on already deployed systems.
@@
(Q4) For which type of transactions do we use which solution? Most LH projects are one-fit-all solutions affecting all transaction types journals/pvldb/ArmbrustDPXZ0YM20 ; delta4_multi_table ; project_nessie_web ; ducklake ; xtable_web . This contradicts a modular solution, providing individual choice between guarantees. However, this limits its simplicity, as such solutions must offer modular building blocks.
3.3 Previous Approaches
Regarding the missing features we identified in Subsection 3.1, we found four relevant designs: External locks journals/pvldb/ArmbrustDPXZ0YM20 , Project Nessie project_nessie_web , Delta Lake 4.0 commit owners delta4_multi_table , and DuckLake ducklake , each representing current approaches.
External Locks journals/pvldb/ArmbrustDPXZ0YM20 follow a locking mechanism (Q1) to sequentialize LH transactions on a table. It is established via an external service like Zookeeper (Q2) and works with the existing OTF structure (Q3). However, due to its table-level granularity, it provides a solution that must be pushed to all (write) transactions (Q4). However, while solving the missing feature (iii), it completely removes concurrency.
Project Nessie project_nessie_web follows similar decisions for Q2 – Q4, but chose a branching mechanism (Q1). It establishes git-like logic for Iceberg, but currently does not allow merging different branches. However, this is essential for the missing feature (ii).
The Delta Lake Commit Owner delta4_multi_table combines versioning and branching (Q1) to provide a central commit via an external, proxy-like, service (Q2). In addition to their external service, they adjust the OTF to fit their design tightly (Q3). While it effectively improves Delta Lake, it also causes compatibility issues with existing implementations. The announced backfilling feature tries to compensate for that, but causes more complexity and delays delta4_multi_table . Lastly, as this approach centralizes commits, the service must be used for all transactions (Q4).
DuckLake ducklake implements multi-table and multi-query transactions by separating the metadata layer from the object store. It manages the metadata of the LH within DuckDB to provide a versioning-based (Q1) solution externally from the object store (Q2). Thus, DuckLake splits the LH design into two distinct parts (Q3), which forces all transactions to go through their structure in DuckDB (Q4). However, we believe such solutions go against the original concept of LHs as they rely on external services, hide the object store behind an external bottleneck, and might reintroduce vendor lock-in.
4 LakeVilla
Figure 2: Overview of all LakeVilla (LV) features and how they integrate into the original OTF (draw.io)
@@ 4 LakeVilla
LakeVilla (LV) is our versioning-based (Q1), built-in (Q2), adaptive (Q3), and modular (Q4) OTF extension providing the missing features (i) – (iii). LV defines three feature sets: LV[R] for (i) Recovery, LV[CT] for (ii) complex transactions, and LV[I] for (iii) isolation. Our design can seamlessly integrate into OTFs such as Apache Iceberg and Delta Lake without reformatting or reconfiguring the whole system.
The Architecture of LV. LV defines modules compatible with OTF operations. Its design does not impact the baseline structure to allow concurrent access to the LH with legacy readers or even other LV combinations. Figure 2 summarizes the typical OTF structure and which modules each LV feature adds. LV[R] and LV[CT] add ’marker placement’ to extend the ’Local version assignment’ of the OTF. Further, LV[R] incorporates a transaction sublog system, and LV[CT] defines concurrent conflict detection across tables. Lastly, LV[I] implements a global layer that assigns (’Global Version Assignment’) and validates versions and changes (’Global Validation’) to guarantee isolation. This section describes all LV features and their modules.
4.1 LakeVilla[Recovery] (LV[R])
@@
Motivation: Why OTF transactions abort. When operating with concurrent transactions, the key challenge to prevent aborts is adapting to changes in the table versions of the metadata layer. By default, OTFs apply optimistic concurrency control that checks for conflicts at the commit of a transaction ib_reliability_docs ; delta_doc ; conf/cidr/Zaharia0XA21 . This approach assumes that most operations do not conflict. However, OTFs also provide table-level granularity, meaning that all concurrent write operations on the same table will conflict without fine-grained conflict management. Hence, we propose LV[R] to provide recovery (i). This feature combines two modules we designed for OTFs following our decisions for Q1–Q4: Our marker placement (module #1) reserves snapshot IDs at a transaction’s beginning seamlessly and publishes a transaction’s changes atomically. To adapt to concurrent changes and faults, we propose our transaction sublogs (module #2).
@@
Figure 3: Module #1: Markers (green) reserve snapshot versions in the metadata layers (draw.io)
@@
Module #1: Marker Placement. We propose file-based markers within OTF metadata layers to prevent concurrent OTF transactions from claiming snapshot versions. We ensure their uniqueness and atomicity by using Put-If-Absent or conditional writes to the object store already used in OTFs aws_conditional_writes ; journals/pvldb/ArmbrustDPXZ0YM20 . Additionally, we implement continuous checks for object stores without atomic requests, accessing each marker’s HEAD to detect potential overwrites. If a transaction fails to claim a version or its marker was deleted by someone else, it aborts or claims a new marker. Markers are temporary metadata objects that must be deleted or overwritten.
Markers in OTFs. In a Delta Lake table, a marker is appended to the log directory. In our example table Bank x, shown in Figure 3, reserves log (3) and its snapshot (3). The executing transaction, consisting of check, withdraw, and deposit operations, can now fully execute and publish all its changes into Log (3) without snapshot (3) being claimed by anybody else. This effectively implements a write lock books/aw/BernsteinHG87 ; journals/pvldb/BernsteinRWY11 ; conf/ipps/IwabuchiSPEGM16 ; journals/pvldb/LiuHMCLSSSAA20 into OTFs. As most OTFs define a synchronization point in their metadata layers journals/pvldb/ArmbrustDPXZ0YM20 ; hudi_web ; iceberg_specs , their marker integration is analogous to the one presented here: For example, in Iceberg, we place markers in the area of manifest lists and link to the snapshot.
@@
Marker Compatibility with Other Clients. Following our design principle of modularity (Q4), we propose that markers not enforce the same guarantees or new techniques on all transactions. Hence, our markers must fulfill their purpose while still being compatible with the original OTF protocols. LV’s markers use the extensible open-source formats (e. g., JSON) of the original logs/manifest and appear empty for the original protocols. At the same time, they are detectable for our marker-based clients with the bookkeeping data contained. This data consists of the transaction ID that created the marker and additional information required for other LV modules (e. g., dependent transactions, time of creation, …).
@@
Marker Commit. Nevertheless, commits using markers should be atomically visible to all clients. So, we designed a commit phase that replaces these markers noticeable to legacy readers: LV collects and summarizes all relevant information according to the OTF’s required instructions during a commit (see Delta Lake in section 2). If needed, it builds underlying structures, such as Iceberg’s manifest lists, and finally overwrites the placed marker with all the necessary instructions. This action effectively makes the whole transaction result of a table available even for the original OTF protocols. In case of an abort, our client overwrites a marker with a previous manifest list or an empty log entry to not disturb the original protocols.